OpenStudy (anonymous):

The function f is defined by f(x) = x2 − 4x + 7 for x > 2. (i) Express f(x) in the form (x − a)2 + b and hence state the range of f. [3] (ii) Obtain an expression for f−1(x) and state the domain of f−1. [3] The function g is defined by g(x) = x − 2 for x > 2. The function h is such that f = hg and the domain of h is x > 0. (iii) Obtain an expression for h(x). [1]

OpenStudy (mathstudent55):

\(f(x) = x^2 − 4x + 7\) \(f(x) = x^2 − 4x ~~~~~~~~~~~~~+ 7\) You need to put in the blank the term that completes the square of \(x^2 - 4x\). Then you need to subtract from the 7 at the end of the line.

OpenStudy (mathstudent55):

Do you know how to complete a square?

OpenStudy (anonymous):

Yeah. Actually it's part 3 I don't know how to do. Could u help with that bit please.?

OpenStudy (mathstudent55):

f = hg h = f/g \(h(x) = \dfrac{f(x)}{g(x)} \)

OpenStudy (anonymous):

Thanks